The average of 11 numbers is 21 . The average of seven of those numbers is 25. What is the average of the remaining four numbers?
 
Please show me step by step how to solve this problem
Thanks

Answer by Alan3354(69443)

The average of 11 numbers is 21 . The average of seven of those numbers is 25. What is the average of the remaining four numbers?
-----------------
average = sum/(# of elements)
--> sum = n*average
sum1 = 11*21 = 231
sum2 = 7*25 = 175
------------------- Subtract
4 numbers total 56
Their avg = 56/4 = 14